You purchased shares of a mutual fund at an offering price of $95 per share at the beginning of the year and paid a front–end load of 4.00%. If the securities in which the fund invested increased in value by 13.25% during the year, and the fund’s expense ratio was 1.50%, what is your rate of return if you sold the fund at the end of the year? Enter your answer rounded to two decimal places.

The price of a large cap equity ETF is $45 at the beginning of the year and $54 at the end of the year. If the ETF paid a $2.25 cash dividend, what is the holding period return for the year? Enter your answer rounded to two decimal places.  

The current market price for ABC is $77 per share. Initial margin is 50%, maintenance margin is 35% and there is no margin interest. ABC pays annual cash dividends of $3.95 per share. You believe the stock price will decrease over the next year and wish to sell short using margin. Suppose you are correct and the stock falls to $62 per share at the end of the year. What is your percentage return on equity for this trade?
Answer #1
1
Purchase price of Mutual Fund $95.00
Front end load @ 4% $3.80
Total Investment $98.80
Sale price of fund $107.59 =95*(1+13.25%)
-Fund's Expense -$1.43 =-95*1.5%
Total Receipt at year end $106.16
Rate of Return
=(106.16-98.8)/98.8
7.45%
2
Total Investment $45.00
Price at the end of the year $54.00
Dividend received $2.25
Total Receipts $56.25
Holding Period return
=(56.25-45)
$11.25
Holding Period Rate of return
=(56.25-45)/45
25.00%
3
Initial Margin Investment @ 50% $38.50 =77/2
Profit on short sell $15.00 =77-62
-Dividend paid -$3.95
Net Gain $11.05
% Return on Trade
=(11.05/38.5)
28.70%